# How Does Message Passing Improve Collaborative Filtering?

### v1

@@ Abstract.
Collaborative filtering (CF) has exhibited prominent results for recommender systems and been broadly utilized for real-world applications. A branch of research enhances CF methods by message passing used in graph neural networks, due to its strong capabilities of extracting knowledge from graph-structured data, like user-item bipartite graphs that naturally exist in CF. They assume that message passing helps CF methods in a manner akin to its benefits for graph-based learning tasks in general (e.g., node classification). However, even though message passing empirically improves CF, whether or not this assumption is correct still needs verification. To address this gap, we formally investigate why message passing helps CF from multiple perspectives and show that many assumptions made by previous works are not entirely accurate. With our curated ablation studies and theoretical analyses, we discover that (i) message passing improves the CF performance primarily by additional representations passed from neighbors during the forward pass instead of additional gradient updates to neighbor representations during the model back-propagation and (ii) message passing usually helps low-degree nodes more than high-degree nodes. Utilizing these novel findings, we present Test-time Aggregation for Collaborative Filtering , namely TAG-CF, a test-time augmentation framework that only conducts message passing once at inference time. The key novelty of TAG-CF is that it effectively utilizes graph knowledge while circumventing most of notorious computational overheads of message passing. Besides, TAG-CF is extremely versatile can be used as a plug-and-play module to enhance representations trained by different CF supervision signals. Evaluated on six datasets (i.e., five academic benchmarks and one real-world industrial dataset), TAG-CF consistently improves the recommendation performance of CF methods without graph by up to 39.2% on cold users and 31.7% on all users, with little to no extra computational overheads. Furthermore, compared with trending graph-enhanced CF methods, TAG-CF delivers comparable or even better performance with less than 1% of their total training time. Beside the promising cost-effectiveness, we show that test-time aggregation in TAG-CF improves recommendation performance in similar ways as the training-time aggregation does, demonstrating the legitimacy of our findings w.r.t. how does message passing improves CF. *Work done during the internship at Snap Inc..
Recommender Systems, Collaborative Filtering, Test-time Augmentation, Efficient Machine Learning Algorithms
1. Introduction
@@
Recommender systems are essential in improving users’ experiences on web services, such as product recommendations (wang2021dcn, ; schafer1999recommender, ), video recommendations (gomez2015netflix, ; van2013deep, ), friend suggestions (sankar2021graph, ; ying2018graph, ), etc. In particular, recommender systems based on collaborative filtering (CF) have shown superior performance (rendle2009bpr, ; koren2021advances, ; chen2023bias, ). CF methods use preferences for items by users to predict additional topics or products a user might like (su2009survey, ). These methods typically learn a unique representation for each user/item and an item is recommended to a user according to the similarity of their representations (he2017neural, ; wang2015collaborative, ).
@@
One popular line of research explores Graph Neural Networks (GNNs) for CF, exhibiting improved results compared with CF frameworks without the utilization of graphs (wu2022graph, ; he2020lightgcn, ; wang2019neural, ; yu2022graph, ; cai2023lightgcl, ; wang2020disentangled, ). The key mechanism behind GNNs is message passing, where each node aggregates information from its neighbors in the graph, and information from neighbors that are multiple hops away can be acquired by stacked message passing layers (kipf2016semi, ; velivckovic2017graph, ; hamilton2017inductive, ). During the model training, traditional CF methods directly fetch user/item representations of an observed interaction (e.g., purchase, friending, click, etc.) and enforce their pair-wise similarity (rendle2009bpr, ). Graph-enhanced CF methods extend this scheme by conducting stacked message passing layers over the user-item bipartite graph, and harnessing the resulting user and item representations to calculate a pair-wise affinity.
@@
A recent study (he2020lightgcn, ) shows that removing several significant components of the message passing layer (e.g., learnable transformation parameters) greatly enhances GNNs’ performance for CF. Its proposed method (LightGCN) achieves promising performance by linearly aggregating neighbor representations without any transformation, and it has been used as the de facto backbone model for later graph-based methods due to its simple and effective design (cai2023lightgcl, ; yu2022graph, ; wu2021self, ). However, this observation contradicts GNN architectures for classic graph learning tasks, where GNNs without these components severely under-perform (oloulade2021graph, ; wang2021bag, ). Additionally, existing research (he2020lightgcn, ; wang2019neural, ) assumes that the contribution of message passing for CF is similar to that for graph learning tasks in general (e.g., node classification or link prediction) - they posit that node representations are progressively refined by their neighbor information and the performance gain is positively proportional to the neighborhood density as measured in node degrees (tang2020investigating, ). However, according to our empirical studies in Section 3.2, message passing in CF improves low-degree users more than high-degree users, which also contradicts GNNs’ behaviors for classic tasks (tang2020investigating, ; hu2022tuneup, ). In light of these inconsistencies between the behaviors of message passing for CF and classic graph learning tasks, we ask:
@@
What role does message passing really
play for collaborative filtering?
@@
In this work, we investigate contributions brought by message passing for CF from two perspectives. Firstly, we unroll the formulation of message passing layer and show that its performance improvement could either come from additional representations passed from neighbors during the forward pass or accompanying gradient updates to neighbor representations during the model back-propagation. With rigorously designed ablation studies, we empirically demonstrate that gains brought by the forward pass dominate those by the back-propagation. Furthermore, we analyze the performance distribution w.r.t. the user degree (i.e., the number of interactions per user) with or without message passing and discover that the message passing in CF improves low-degree users more compared to high-degree users. For the first time, we connect this phenomenon to Laplacian matrix learning (zhu2021interpreting, ; dong2019learning, ; dong2016learning, ), and theoretically show that popular supervision signals (rendle2009bpr, ; wang2022towards, ) for CF inadvertently conduct message passing in the back-propagation even without treating the input data as a graph. Hence, when message passing is applied, high-degree users demonstrate limited improvement, as the benefit of message passing for high degree nodes has already been captured by the supervision signal.
@@
With the above takeaways, we present Test-time Aggregation for Collaborative Filtering , namely TAG-CF. Specifically, unlike other graph CF methods, TAG-CF does not require any message passing during training. Instead, it is a test-time augmentation framework that only conducts a single message-passing step at inference time, and effectively enhances representations inferred from different CF supervision signals. The test-time aggregation is inspired by our first perspective that, within total performance gains brought by message passing, gains from the forward pass dominate those brought by the additional back-propagation. Applying message passing only at test time avoids repetitive queries (i.e., once per node and epoch) for representations of surrounding neighbors, which grow exponentially as the number of layers increases. Moreover, following our second perspective that message passing helps low-degree nodes more in CF, we further offload the cost of TAG-CF by applying the one-time message passing only to low-degree nodes. In short, we summarize our contributions as:
@@
2. Preliminaries and Related Work
Collaborative Filtering. Given a set of users, a set of items, and interactions between users and items, collaborative filtering (CF) methods aim at learning a unique representation for each user and item, such that user and item representations can reconstruct all observable interactions (rendle2009bpr, ; wang2022towards, ; koren2009matrix, ). CF methods based on matrix factorization directly utilize the inner product between a pair of user and item representations to infer the existence of their interaction (koren2009matrix, ; rendle2009bpr, ). Further leveraging nonlinearities, CF methods based on neural predictors use multi-layer feed-forward neural networks that take user and item representations as inputs and output prediction results (he2017neural, ; zhang2019deep, ). Let $\mathcal{U}$ and $\mathcal{I}$ denote the user set and item set respectively, with user $u_{i}\in\mathcal{U}$ associated with an embedding $\mathbf{u}_{i}\in\mathbb{R}^{d}$ and item $i_{i}\in\mathcal{U}$ associated with $\mathbf{i}_{i}\in\mathbb{R}^{d}$, the similarity $s_{ij}$ between user $u_{i}$ and item $i_{j}$ is formulated as $s_{ij}=\hat{\mathbf{u}}_{i}^{\intercal}\cdot\hat{\mathbf{i}}_{j}$.
Graph Neural Networks. Graph neural networks (GNNs) are powerful learning frameworks to extract representative information from graphs (kipf2016semi, ; velivckovic2017graph, ; hamilton2017inductive, ; xu2018representation, ; ju2023let, ; fan2022heterogeneous, ; ju2023let, ). They aim to map each input node into low-dimensional vectors, which can be utilized to conduct either graph-level (xu2018powerful, ) or node-level tasks (kipf2016semi, ). Most GNNs explore layer-wise message passing (gilmer2017neural, ), where each node iteratively extracts information from its first-order neighbors, and information from multi-hop neighbors can be captured by stacked layers. Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and node features $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$, graph convolution in GCN (kipf2016semi, ) at $k$-th layer is formulated as:
(1) $\mathbf{h}_{i}^{(k+1)}=\sigma(\sum_{j\in\mathcal{N}(i)\cup{i}}\frac{1}{\sqrt{|N(i)|}\sqrt{|N(j)|}}\mathbf{h}_{j}^{(k)}\cdot\mathbf{W}^{(k)}),$
@@ 2. Preliminaries and Related Work
Recent works (ma2021homophily, ; ma2021unified, ) have shown that GNNs make predictions based on the distribution of node neighborhoods. Moreover, GNNs’ performance improvement for high-degree nodes is typically better than for low-degree nodes (tang2020investigating, ; hu2022tuneup, ; ju2024graphpatcher, ; guo2024node, ). They posit that node representations are progressively refined by their neighbor information and the performance gain is positively proportional to the neighborhood density as measured in node degrees.
Message Passing for Collaborative Filtering. Recent research tends to apply the message passing scheme in GNNs to CF (he2020lightgcn, ; wang2019neural, ; pal2020pinnersage, ; gao2022graph, ; shi2020heterogeneous, ; xia2021graph, ). In CF, they mostly conduct message passing between user-item bipartite graphs and utilize the resultant representations to calculate user-item similarities. For instance, NGCF (wang2019neural, ) directly migrates the message passing scheme in GNNs (similar to Equation 1) and applies it to bipartite graphs in CF. LightGCN (he2020lightgcn, ) simplifies NGCF (wang2019neural, ) by removing certain components (i.e., the self-loop, learning parameters for graph convolution, and activation functions) and further improves the recommendation performance compared with NGCF. The simplified parameter-less message passing in LightGCN can be expressed as:
(2) $\mathbf{u}_{i}^{(k)}=\sum_{i_{j}\in N(u_{i})}\frac{1}{\sqrt{|N(u_{i})|}\sqrt{|N(i_{j})|}}\mathbf{i}_{j}^{(k-1)},$
(3) $\mathbf{i}_{i}^{(k)}=\sum_{u_{j}\in N(i_{i})}\frac{1}{\sqrt{|N(i_{i})|}\sqrt{|N(u_{j})|}}\mathbf{u}_{j}^{(k-1)},$
@@
(4) $\hat{\mathbf{u}}_{i}=\frac{1}{K+1}\sum_{k=0}^{K}\mathbf{u}_{i}^{(k)},\;\;\hat{\mathbf{i}}_{i}=\frac{1}{K+1}\sum_{k=0}^{K}\mathbf{i}_{i}^{(k)},\;\;s_{ij}=\hat{\mathbf{u}}_{i}^{\intercal}\cdot\hat{\mathbf{i}}_{j}.$
According to results reported in LightGCN and NGCF (he2020lightgcn, ; wang2019neural, ; chang2021sequential, ; gao2023survey, ) and empirical studies we provide in this work (i.e., Table 2 and Table 5), incorporating message passing to CF methods without graphs (i.e., matrix factorization methods (rendle2009bpr, ; he2017neural, )) can improve the recommendation performance by up to 20%. Utilizing LightGCN as the backbone model, later works try to further improve the performance by incorporating self-supervised learning signals (lin2022improving, ; yu2022graph, ; cai2023lightgcl, ; yu2021self, ; wei2021contrastive, ; ju2022multi, ). Graph-based CF methods assume that the contribution of message passing for CF is similar to that for graph learning tasks in general (e.g., node classification or link prediction). However, whether or not this assumption is correct still needs verification, even though message passing empirically improves CF. There also exists a branch of research that aims at accelerating or simplifying message passing in CF by adding graph-based regularization terms during the training (shen2021powerful, ; mao2021ultragcn, ; peng2022svd, ; xia2023graph, ). While promising, they still repetitively query representations of adjacent nodes during the training.
3. How Does Message Passing Improve Collaborative Filtering?
In this section, we demonstrate the reason behind why message passing helps collaborative filtering from two major perspectives: Firstly, we focus on inductive biases brought by the message passing explored in LightGCN, the de facto backbone model for graph-based CF methods. Secondly, we consider the performance improvement on different node subgroups w.r.t. the node degree with and without message passing.
3.1. Neighbor Information vs. Accompanying Gradients from Message Passing
@@ 3. How Does Message Passing Improve Collaborative Filtering?
$s_{ij}=\Big{(}\mathbf{u}_{i}+\sum_{i_{n}\in N(u_{i})}\frac{1}{\sqrt{|N(u_{i})|}\sqrt{|N(i_{n})|}}\mathbf{i}_{n}\Big{)}^{\intercal}$
$\;\;\;\;\;\cdot\Big{(}\mathbf{i}_{j}+\sum_{u_{n}\in N(i_{j})}\frac{1}{\sqrt{|N(i_{j})|}\sqrt{|N(u_{n})|}}\mathbf{u}_{n}\Big{)}$
$\;\;\;=\;\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}+\sum_{u_{n}\in N(i_{j})}\frac{1}{\sqrt{|N(i_{j})|}\sqrt{|N(u_{n})|}}\mathbf{u}_{i}^{\intercal}\cdot\mathbf{u}_{n}$
(5) $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\sum_{i_{n}\in N(u_{i})}\frac{1}{\sqrt{|N(u_{i})|}\sqrt{|N(i_{n})|}}\mathbf{i}_{n}^{\intercal}\cdot\mathbf{i}_{j}$
$+\sum_{i_{n}\in N(u_{i})}\sum_{u_{n}\in N(i_{j})}\frac{1}{\sqrt{|N(u_{i})|}\sqrt{|N(i_{n})|}\sqrt{|N(i_{j})|}\sqrt{|N(u_{n})|}}\mathbf{i}_{n}^{\intercal}\cdot\mathbf{u}_{n}.$
With derived similarities between user-item pairs, their corresponding representations can be updated by objectives (e.g., BPR (rendle2009bpr, ) and DirectAU (wang2022towards, )) that enforce the pair-wise similarity between representations of user-item pairs in the training data.
Table 1. Performance of LightGCN variants.
@@
CF methods without the utilization of graphs directly calculate the similarity between a user and an item with their own representations (i.e., $s_{ij}=\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}$), which aligns with the first term in Equation 5. Compared to the formulation in Equation 5, we can see that three additional similarity terms are introduced as inductive biases: similarities between users who purchase the same item (i.e., $\mathbf{u}_{i}^{\intercal}\cdot\mathbf{u}_{n}$), between items that share the same buyer (i.e., $\mathbf{i}_{n}^{\intercal}\cdot\mathbf{i}_{j}$), and between neighbors of an observed interaction (i.e., $\mathbf{i}_{n}^{\intercal}\cdot\mathbf{u}_{n}$). With these three additional terms from message passing, we reason that the performance improvement brought by message passing to CF methods without graph could come from (i) additional neighbor representations during the forward pass (i.e., numerical values of three extra terms in Equation 5), or (ii) accompanying gradient updates to neighbors during the back-propagation.
To investigate the origin of the performance improvement brought by message passing, we designed two variants of LightGCN. The first one (LightGCNw/o neigh. ${}_{\text{w/o neigh. info}}$) shares the same forward and backward procedures as LightGCN during the training but does not conduct message passing during the test time. In this variant, additional gradients brought by message passing are maintained as part of the resulting model, but information from neighbors are ablated. In the second variant (LightGCNw/o ${}_{\text{w/o grad.}}$), the model shares the same forward pass but drops gradients from these three additional terms during the backward propagation. Besides these two variants, we also experiment on LightGCN without message passing, denoted as LightGCNw/o ${}_{\text{w/o both}}$, a matrix factorization model with the same supervision signal (i.e., BPR loss). Implementation details w.r.t. this experiment are in Appendix C.
Figure 1. Performances of LightGCN and Matrix Factorization w.r.t. the user degree across datasets. The performance improvement brought by message passing decreases as the user degree goes up.
@@
From Table 1, we observe that the performance of all variants is downgraded compared with LightGCN, with the most significant degradation on LightGCNw/o neigh. ${}_{\text{w/o neigh. info}}$. This phenomenon indicates that (i) both additional representations passed from neighbors during the forward pass and accompanying gradient updates to neighbors during the back-propagation help the recommendation performance, and (ii) within total performance gains brought by message passing, gains from the forward pass dominate those brought by the back-propagation. Comparing LightGCN with LightGCNw/o ${}_{\text{w/o grad.}}$, we notice that the incorporation of gradient updates brought by message passing is relatively incremental (i.e., $\sim$2%). However, to facilitate these additional gradient updates for slightly better performance, LightGCN is required to conduct message passing at each batch, which brings tremendous additional overheads.
@@
3.2. Message Passing in CF Helps Low-degree Users More Compared with High-degrees

### v2

@@ Abstract
Collaborative filtering (CF) has exhibited prominent results for recommender systems and been broadly utilized for real-world applications. A branch of research enhances CF methods by message passing (MP) used in graph neural networks, due to its strong capabilities of extracting knowledge from graph-structured data, like user-item bipartite graphs that naturally exist in CF. They assume that MP helps CF methods in a manner akin to its benefits for graph-based learning tasks in general (e.g., node classification). However, even though MP empirically improves CF, whether or not this assumption is correct still needs verification. To address this gap, we formally investigate why MP helps CF from multiple perspectives and show that many assumptions made by previous works are not entirely accurate. With our curated ablation studies and theoretical analyses, we discover that (i) MP improves the CF performance primarily by additional representations passed from neighbors during the forward pass instead of additional gradient updates to neighbor representations during the model back-propagation and (ii) MP usually helps low-degree nodes more than high-degree nodes. Utilizing these novel findings, we present Test-time Aggregation for Collaborative Filtering , namely TAG-CF, a test-time augmentation framework that only conducts MP once at inference time. The key novelty of TAG-CF is that it effectively utilizes graph knowledge while circumventing most of notorious computational overheads of MP. Besides, TAG-CF is extremely versatile can be used as a plug-and-play module to enhance representations trained by different CF supervision signals. Evaluated on six datasets (i.e., five academic benchmarks and one real-world industrial dataset), TAG-CF consistently improves the recommendation performance of CF methods without graph by up to 39.2% on cold users and 31.7% on all users, with little to no extra computational overheads. Furthermore, compared with trending graph-enhanced CF methods, TAG-CF delivers comparable or even better performance with less than 1% of their total training times. Our code is publicly available at https://github.com/snap-research/Test-time-Aggregation-for-CF.
1 Introduction
@@
Recommender systems are essential in improving users’ experiences on web services, such as product recommendations [59, 47], video recommendations [15, 55], friend suggestions [46, 72], etc. In particular, recommender systems based on collaborative filtering (CF) have shown superior performance [45, 33, 4]. CF methods use preferences for items by users to predict additional topics or products a user might like [51]. These methods typically learn a unique representation for each user/item and an item is recommended to a user according to their representation similarities [21, 58].
@@
One popular line of research explores Graph Neural Networks (GNNs) for CF, exhibiting improved results compared with CF frameworks without the utilization of graphs [66, 20, 60, 74, 1, 61, 80]. The key mechanism behind GNNs is message passing (MP), where each node aggregates information from its neighbors in the graph, and information from neighbors that are multiple hops away can be acquired by stacked MP layers [30, 56, 18]. During the model training, traditional CF methods directly fetch user/item representations of an observed interaction (e.g., purchase, friending, click, etc.) and enforce their pair-wise similarity [45]. Graph-enhanced CF methods extend this scheme by conducting stacked MP layers over the user-item bipartite graph, and harnessing the resulting user and item representations to calculate a pair-wise affinity.
@@
A recent study [20] shows that removing several key components of the MP layer (e.g., learnable transformation parameters) greatly enhances GNNs’ performance for CF. Its proposed method (LightGCN) achieves promising performance by linearly aggregating neighbor representations and has been used as the de facto backbone model for later works due to its simple and effective design [1, 74, 65]. However, this observation contradicts GNN architectures for classic graph learning tasks, where GNNs without these components severely under-perform [41, 62]. Additionally, existing research [20, 60] assumes that the contribution of MP for CF is similar to that for graph learning tasks in general (e.g., node classification or link prediction) - they posit that node representations are progressively refined by their neighbor information and the performance gain is positively proportional to the neighborhood density as measured in node degrees [54]. However, according to our empirical studies in Section 3.2, MP in CF improves low-degree nodes more than high-degree nodes, which also contradicts GNNs’ behaviors for classic tasks [54, 23]. In light of these inconsistencies, we ask:
@@
What role does message passing really play for collaborative filtering?
@@
In this work, we investigate contributions brought by MP for CF from two perspectives. Firstly, we unroll the formulation of MP layer and show that its performance improvement could either come from additional representations passed from neighbors during the forward pass or accompanying gradient updates to neighbor representations during the back-propagation. With rigorously designed ablation studies, we empirically demonstrate that gains brought by the forward pass dominate those by the back-propagation. Furthermore, we analyze the performance distribution w.r.t. the user degree (i.e., the number of interactions per user) with or without message passing and discover that the message passing in CF improves low-degree users more compared to high-degree users. For the first time, we connect this phenomenon to Laplacian matrix learning [82, 10, 9], and theoretically show that popular supervision signals [45, 57] for CF inadvertently conduct message passing in the back-propagation even without treating the input data as a graph. Hence, when message passing is applied, high-degree users demonstrate limited improvement, as the benefit of message passing for high degree nodes has already been captured by the supervision signal.
@@
With the above takeaways, we present Test-time Aggregation for Collaborative Filtering , namely TAG-CF. Specifically, unlike other graph CF methods, TAG-CF does not require any message passing during training. Instead, it is a test-time augmentation framework that only conducts a single message-passing step at inference time, and effectively enhances representations inferred from different CF supervision signals. The test-time design is inspired by our first perspective that, within total performance gains brought by message passing, gains from the forward pass dominate those brought by the backward pass. Applying message passing only at test time avoids repetitive queries (i.e., once per node and epoch) for representations of surrounding nodes, which grow exponentially as the number of layers increases. Moreover, following our second perspective that message passing helps low-degree nodes more in CF, we further offload the cost of TAG-CF by applying the one-time message passing only to low-degree nodes. We summarize our contributions as:
@@
2 Preliminaries and Related Work
Collaborative Filtering. Given a set of users, a set of items, and interactions between users and items, collaborative filtering (CF) methods aim at learning a unique representation for each user and item, such that user and item representations can reconstruct all observable interactions [45, 57, 32]. CF methods based on matrix factorization directly utilize the inner product between a pair of user and item representations to infer the existence of their interaction [32, 45]. CF methods based on neural predictors use multi-layer feed-forward neural networks that take user and item representations as inputs and output prediction results [21, 77]. Let $\mathcal{U}$ and $\mathcal{I}$ denote the user set and item set respectively, with user $u_{i}\in\mathcal{U}$ associated with an embedding $\mathbf{u}_{i}\in\mathbb{R}^{d}$ and item $i_{i}\in\mathcal{U}$ associated with $\mathbf{i}_{i}\in\mathbb{R}^{d}$, the similarity $s_{ij}$ between user $u_{i}$ and item $i_{j}$ is formulated as $s_{ij}=\hat{\mathbf{u}}_{i}^{\intercal}\cdot\hat{\mathbf{i}}_{j}$.
Graph Neural Networks. Graph neural networks (GNNs) are powerful learning frameworks to extract representative information from graphs [30, 56, 18, 70, 11, 26], with numerous applications in large-scale ranking and forecasting tasks [53, 52, 8, 49]. They aim to map each input node into low-dimensional vectors, which can be utilized to conduct either graph-level [69] or node-level tasks [30]. Most GNNs explore layer-wise message passing [14], where each node iteratively extracts information from its first-order neighbors, and information from multi-hop neighbors can be captured by stacked layers. Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and node features $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$, graph convolution in GCN [30] at $k$-th layer is formulated as:
$\mathbf{h}_{i}^{(k+1)}=\sigma(\sum_{j\in\mathcal{N}(i)\cup{i}}\frac{1}{\sqrt{|N(i)|}\sqrt{|N(j)|}}\mathbf{h}_{j}^{(k)}\cdot\mathbf{W}^{(k)}),$ (1)
@@ 2 Preliminaries and Related Work
Recent works [37, 38] have shown that GNNs make predictions based on the distribution of node neighborhoods. Moreover, GNNs’ performance improvement for high-degree nodes is typically better than for low-degree nodes [54, 23, 28, 17, 63]. They posit that node representations are progressively refined by their neighbor information and the performance gain is positively proportional to the neighborhood density as measured in node degrees. As we explore test-time augmentation in this work, it is worth noting that there also exist a group of relevant works that explore data augmentation techniques to enhance the GNN performance [28, 79, 78, 25, 24].
Message Passing for Collaborative Filtering. Recent research tends to apply the message passing scheme in GNNs to CF [20, 60, 42, 12, 50, 68, 31, 34]. In CF, they mostly conduct message passing between user-item bipartite graphs and utilize the resultant representations to calculate user-item similarities. For instance, NGCF [60] directly migrates the message passing scheme in GNNs (similar to Equation 1) and applies it to bipartite graphs in CF. LightGCN [20] simplifies NGCF [60] by removing certain components (i.e., the self-loop, learning parameters for graph convolution, and activation functions) and further improves the recommendation performance compared with NGCF. The simplified parameter-less message passing in LightGCN can be expressed as:
$\mathbf{u}_{i}^{(k)}=\sum_{i_{j}\in N(u_{i})}\frac{1}{\sqrt{|N(u_{i})|}\sqrt{|N(i_{j})|}}\mathbf{i}_{j}^{(k-1)},\mathbf{i}_{i}^{(k)}=\sum_{u_{j}\in N(i_{i})}\frac{1}{\sqrt{|N(i_{i})|}\sqrt{|N(u_{j})|}}\mathbf{u}_{j}^{(k-1)},$ (2)
@@
$\hat{\mathbf{u}}_{i}=\frac{1}{K+1}\sum_{k=0}^{K}\mathbf{u}_{i}^{(k)},\;\;\hat{\mathbf{i}}_{i}=\frac{1}{K+1}\sum_{k=0}^{K}\mathbf{i}_{i}^{(k)},\;\;s_{ij}=\hat{\mathbf{u}}_{i}^{\intercal}\cdot\hat{\mathbf{i}}_{j}.$ (3)
According to results reported in LightGCN and NGCF [20, 60, 2, 13] and empirical studies we provide in this work (i.e., Table 2 and Section 5.2), incorporating message passing to CF methods without graphs (i.e., matrix factorization methods [45, 21]) can improve the recommendation performance by up to 20%. Utilizing LightGCN as the backbone model, later works try to further improve the performance by incorporating self-supervised learning signals [35, 74, 1, 73, 64, 27]. Graph-based CF methods assume that the contribution of message passing for CF is similar to that for graph learning tasks in general (e.g., node classification or link prediction). However, whether or not this assumption is correct still needs verification, even though message passing empirically improves CF. There also exists a branch of research that aims at accelerating or simplifying message passing in CF by adding graph-based regularization terms during the training [48, 39, 44, 67]. While promising, they still repetitively query representations of adjacent nodes during the training.
Efficient Efforts in Matrix Factorization. A branch of research specifically focuses on improving the efficiency of matrix factorization [48, 44, 22, 43, 6]. For instance, GFCF [48] and Turbo-CF [43] explore graph signal processing to linearly convolve the interaction matrix and use the resulted matrix directly for recommendation without training. Furthermore, SVD-GCN [44] and SVD-AE [22] utilize a low rank version of the interaction matrix to further accelerate the convolution efficiency and yet remain the promising performance. Besides, BSPM [6] studies using diffusion process to gradually reconstruct the interaction matrix and achieves promising performance with fast processing. In parallel with these existing efforts, we propose to enhance any existing matrix factorization method through test-time augmentation that harnesses graph-based heuristics.
3 How Does Message Passing Improve Collaborative Filtering?
In this section, we demonstrate why message passing (MP) helps collaborative filtering from two major perspectives: Firstly, we focus on inductive biases brought by the MP explored in LightGCN, the de facto backbone model for graph-based CF methods. Secondly, we consider the performance improvement on different node subgroups w.r.t. the node degree with and without MP.
3.1 Neighbor Information vs. Accompanying Gradients from Message Passing
@@ 3 How Does Message Passing Improve Collaborative Filtering?
$s_{ij}=\Big{(}\mathbf{u}_{i}+\sum_{i_{n}\in N(u_{i})}\frac{1}{\sqrt{|N(u_{i})|}\sqrt{|N(i_{n})|}}\mathbf{i}_{n}\Big{)}^{\intercal}\cdot\Big{(}\mathbf{i}_{j}+\sum_{u_{n}\in N(i_{j})}\frac{1}{\sqrt{|N(i_{j})|}\sqrt{|N(u_{n})|}}\mathbf{u}_{n}\Big{)}$
$\;\;\;=\;\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}+\sum_{u_{n}\in N(i_{j})}\frac{1}{\sqrt{|N(i_{j})|}\sqrt{|N(u_{n})|}}\mathbf{u}_{i}^{\intercal}\cdot\mathbf{u}_{n}\sum_{i_{n}\in N(u_{i})}\frac{1}{\sqrt{|N(u_{i})|}\sqrt{|N(i_{n})|}}\mathbf{i}_{n}^{\intercal}\cdot\mathbf{i}_{j}$
$\;\;\;+\sum_{i_{n}\in N(u_{i})}\sum_{u_{n}\in N(i_{j})}\frac{1}{\sqrt{|N(u_{i})|}\sqrt{|N(i_{n})|}\sqrt{|N(i_{j})|}\sqrt{|N(u_{n})|}}\mathbf{i}_{n}^{\intercal}\cdot\mathbf{u}_{n}.$ (4)
With derived similarities between user-item pairs, their corresponding representations can be updated by objectives (e.g., BPR [45] and DirectAU [57]) that enforce the pair-wise similarity between representations of user-item pairs in the training data.
CF methods without the utilization of graphs directly calculate the similarity between a user and an item with their own representations (i.e., $s_{ij}=\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}$), which aligns with the first term in Equation 4. Compared to the formulation in Equation 4, we can see that three additional similarity terms are introduced as inductive biases: similarities between users who purchase the same item (i.e., $\mathbf{u}_{i}^{\intercal}\cdot\mathbf{u}_{n}$), between items that share the same buyer (i.e., $\mathbf{i}_{n}^{\intercal}\cdot\mathbf{i}_{j}$), and between neighbors of an observed interaction (i.e., $\mathbf{i}_{n}^{\intercal}\cdot\mathbf{u}_{n}$). With these three additional terms from MP, we reason that the performance improvement brought by MP to CF methods without graph could come from (i) additional neighbor representations during the forward pass (i.e., numerical values of three extra terms in Equation 4), or (ii) accompanying gradient updates to neighbors during the back-propagation.
To investigate the origin of the performance improvement brought by MP, we designed two variants of LightGCN. The first one (LightGCNw/o neigh. ${}_{\text{w/o neigh. info}}$) shares the same forward and backward procedures as LightGCN during the training but does not conduct MP during the test time. In this variant, additional gradients brought by MP are maintained as part of the resulting model, but information from neighbors are ablated. In the second variant (LightGCNw/o ${}_{\text{w/o grad.}}$), the model shares the same forward pass but drops gradients from these three additional terms during the backward propagation. Besides these two variants, we also experiment on LightGCN without MP, denoted as LightGCNw/o ${}_{\text{w/o both}}$, a matrix factorization model with the same supervision signal (i.e., BPR loss). Implementation details w.r.t. this experiment are in Appendix C.
Table 1: Performance of LightGCN variants.
@@
From Table 1, we observe that the performance of all variants is downgraded compared with LightGCN, with the most significant degradation on LightGCNw/o neigh. ${}_{\text{w/o neigh. info}}$. This phenomenon indicates that (i) both additional representations passed from neighbors during the forward pass and accompanying gradient updates to neighbors during the back-propagation help the recommendation performance, and (ii) within total performance gains brought by MP, gains from the forward pass dominate those brought by the back-propagation. Comparing LightGCN with LightGCNw/o ${}_{\text{w/o grad.}}$, we notice that the incorporation of gradient updates brought by MP is relatively incremental (i.e., $\sim$2%). However, to facilitate these additional gradient updates for slightly better performance, LightGCN is required to conduct MP at each batch, which brings tremendous additional overheads.
@@
3.2 Message Passing in CF Helps Low-degree Users More Compared with High-degrees